OpenStudy (anonymous):

Help with the 3rd page?

OpenStudy (anonymous):

Can I get some help with the 3rd page in this file? I have completed page one and 2 but confused by page 3.

OpenStudy (ddcamp):

On the 3rd page 1a is a simple maths error. 5 + (12-1)*2 = 5 + 11* 2 = 5+22, NOT 16*2 1b doesn't have any problems other than an incorrect a₁₂ value. Problem 2 isn't an arithmetic sequence, but they used the formula for the sum of an arithmetic sequence. Problem 3 starts at n=0, not n=1 like the formula implies Problem 4 While -2 < 1, |-2| > 1, so it will diverge.

OpenStudy (anonymous):

Thank you so much! I completed page 1 and 2 but I was just having trouble finding the errors with the last 4 problems